OpenStudy (anonymous):

Help Please! MEDAL! Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2 (1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (anonymous):

-5<-3(1)+1 5<-3(1)+3 1<-3(5)+3 5=-3(-1)+3

OpenStudy (anonymous):

the first one is -5<-3+1=-5<-2= correct the second one is 5<-3(1)+3 5<-3+3 5<0 the thrid one is 1<-15+3=1<-12 and the last one is 5<3+3=5<6 so the first and last one are correct.

OpenStudy (anonymous):

then tou must do it for the second equation -5<1+2=-5<3= correct and 5<-1+2=5<-1= incorrect

OpenStudy (anonymous):

so the answer is number 1

OpenStudy (anonymous):

or (1,-5)